"Find the slope of each line"
y= 4/3x -2
what are the steps to solve the solution?

Answer by jim_thompson5910(35256)

y= 4/3x -2 is in the form y = mx+b where m = 4/3. So the slope is 4/3.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------